OpenStudy (haleyelizabeth2017):

The design of a digital box camera maximizes the volume while keeping the sum of the dimensions at 6 inches. If the length must be 1.5 times the height, what should each dimension be? *Hint: Let x represent one of the dimensions, and then define the other dimensions in terms of x.

OpenStudy (haleyelizabeth2017):

@TQKMB this is the question...

OpenStudy (haleyelizabeth2017):

@SolomonZelman :/ so for the first equation wouldn't it be x+1.5x+y=6 or 2.5x+y=6?

OpenStudy (solomonzelman):

"If the length must be 1.5 times the height" \(\large\color{black}{ l=1.5w }\) "the sum of the dimensions at 6 inches" \(\large\color{black}{ l+w +h=6 }\)

OpenStudy (solomonzelman):

\(\large\color{black}{ 1.5w+w +h=6 }\) \(\large\color{black}{ 2.5w+h=6 }\)

OpenStudy (solomonzelman):

and you want to find a maximum value with (will say of a surface) when your dimensions are 2.5 and 6. looks like an optimization problem?

OpenStudy (solomonzelman):

I mean not dimensions 2.5 and 6, but w and h

OpenStudy (solomonzelman):

Is this calculus?

OpenStudy (haleyelizabeth2017):

oops! So sorry....no it is not....it's Honors Algebra 2

OpenStudy (solomonzelman):

well, either way, you want to find all dimensions in terms of 1 variable, and you have a 3D box, that right? or 2D box?

OpenStudy (haleyelizabeth2017):

this section is polynomials and polynomial functions...

OpenStudy (haleyelizabeth2017):

ummm.....I'm not sure....what I put in the question is all they gave me :(

OpenStudy (solomonzelman):

if it is algebra 2, then there are only 2 dimensions, because they don't want to do optimization problem from calc 1, although it is a very typical optimization prob from calc. \(\large\color{black}{ l=1.5w }\) \(\large\color{black}{ l+w=6 }\) \(\large\color{black}{ 1.5w+w=6 }\) \(\large\color{black}{ 2.5w=6 }\) \(\large\color{black}{ w=2.4 }\) \(\large\color{black}{ l=2.4\times 1.5=3.6 }\)

OpenStudy (solomonzelman):

well, I don't understand then, why do are you given that you want the maximum volume if you already have 2 equations given (1. sum of Dimensions is 6, 2. length is 1.5 times greater than the width) ?

OpenStudy (solomonzelman):

see what I am asking ?

OpenStudy (haleyelizabeth2017):

I'm not sure :( this is just super confuzzling

OpenStudy (solomonzelman):

But on the other hand, it can't be an optimization problem with 3 dimensions, because you are just in alg 2.

OpenStudy (solomonzelman):

I would say just 2 dimensions, the way I did it.

OpenStudy (haleyelizabeth2017):

okay...

OpenStudy (haleyelizabeth2017):

Is that all I have to do?

OpenStudy (solomonzelman):

apparently.

OpenStudy (haleyelizabeth2017):

okay thank you so much!

OpenStudy (solomonzelman):

yw